The sum of the ages of three sisters, Carol, Cindy, and Lisa, is 42. Carol, the oldest, is 4 times as old as Lisa, who is the youngest. How old is Cindy? Explain how you know.
Answer by ankor@dixie-net.com(22740)

The sum of the ages of three sisters, Carol, Cindy, and Lisa, is 42.
Carol, the oldest, is 4 times as old as Lisa, who is the youngest.
How old is Cindy? Explain how you know.
:
let x, y, z = the three sisters, x is youngest, z the oldest
:
x + y + z = 42
:
The oldest (z) 4 times the youngest (x)
z = 4x
:
Replace z with 4x in the sum equation
x + y + 4x = 42
y + 5x = 42
y = 42 - 5x
:
Some values for this equation:
x | y
-----
2 |32
4 |22
5 |17
6 |12
7 | 7
8 | 2
:
know y<4x and y>x, only two values for x will satisfy this:
x=5, y=17, z=20
and
x=6, y=12, z=24
:
Two solutions,
Cindy is age 17, or Cindy is age 12